# A Low Duty Cycle Pulsed UV Technique for Spectroscopy of Aluminum Monochloride

### v1

@@
A Low Duty Cycle Pulsed UV Technique for Spectroscopy of Aluminum Monochloride
@@ 1 Introduction
The invention of the laser led to a myriad of applications in science and industry [1]. In fundamental research, the laser enabled applications ranging from precision spectroscopy and tests of fundamental theories [2], cooling and trapping of atoms and molecules [3, 4], to cold controlled chemistry [5, 6, 7]. This type of research often uses narrow-band continuous-wave (CW) lasers to provide a high frequency resolution. Many of these applications also require laser light in the deep ultraviolet (UV) to cover optical transitions in atoms and molecules for carrying out precision spectroscopy and laser cooling [3]. Specific examples include atomic cadmium with a laser cooling transition at 229 nm [8, 9] and atomic mercury at 254 nm [10, 11].
@@
In addition to a specific wavelength in the deep-UV, many of the interesting applications require the use of high UV laser power. For instance, driving dipole-forbidden transitions, as is done for spectroscopy of hydrogen [14, 15, 16, 17], muonium [18] and xenon [19], requires deep-UV laser light at the level of Watts. Another example involves the laser cooling of molecules [20, 4, 21, 22]. Two of the promising candidates, AlCl [23, 24] and AlF [25], have laser cooling transitions at 261 nm and 228 nm, respectively. These molecules potentially provide high capture velocities for magneto-optical traps of up to 30–40 m/s [24]. To achieve this level, though, high intensities are required to saturate the cooling transitions. At the same time, this requirement is in conflict with the desire to use laser beams with large beam cross sections. Here, the aim is to provide a large overlap with a molecular beam to render the capture process as efficient as possible. To address such experimental needs, tremendous progress has been made over the years and various laser systems at wavelengths in the deep-UV below 300 nm with output levels ranging from 50 mW to more than 1 Watt, have been developed [26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37].
@@
A major challenge when using deep-UV lasers is the laser-induced damage that can occur in any of the involved optical components. The intensity threshold, typically designated as laser-induced damage threshold (LIDT), at which significant damage occurs depends on the optical material itself and varies with the wavelength. The mechanisms of these effects have been studied for typical used optical materials, e.g. CaF2 [38, 39] and fused silica [40, 41, 42]. Among the detrimental effects causing damage is the absorption of moisture by materials that are hygroscopic, the breaking of chemical bonds, the depletion of oxygen and the contamination with hydrocarbons upon exposure [43, 44, 45]. Various methods have been explored to revert these effects that focus on keeping a oxygen near the surface of the optics through oxide coatings [45] or the submersion of the optics in an oxygen environment [32]. Moreover, fluoride-coated optics that avoid the need for the presence of oxygen have been shown to be beneficial [46]. Non-linear crystals, which are used for higher-harmonic generation to produce deep-UV light, are often hygroscopic and suffer from UV-induced damage. Extensive studies have been carried out to characterize these damages and to find mitigation techniques for non-linear crystals, such as BaB2O4 (BBO) [47, 48, 49] and CsLiB6O10 (CLBO) [50, 51, 52, 53, 49], which are relevant for this work. Optical fibers, another essential component in today’s experiments involving lasers, have been shown to be able to withstand larger amounts of UV intensities if they are pretreated with hydrogen and irradiate with UV light [54].
@@
While the choice of materials and the developed specific treatments have shown promising results to prepare optical components for operating in the deep-UV regime, a material-independent method to reduce UV-degradation in the first place is to minimize high intensity UV-exposure of an optical material. In some optical setups this is possible by avoiding laser beam foci near or inside optical materials, which consequently reduces the local UV intensity [55]. This technique, however, cannot be used in setups where optical foci are a requirement. This is the case for resonant enhancement cavities that are typically used to efficiently implement the higher harmonic generation of CW laser light by focusing the cavity mode into a non-linear conversion crystal to maximize the electric field inside the crystal [56]. Here, an alternative approach is to reduce the duty cycle of the UV light to a minimum instead. This technique is most naturally applied in pulsed experiments that often exhibit times between experimental repeats where the laser is only idling. However, it can also be used to bridge dead times in continuous experiments.
@@
In this work, we present a method to reduce the duty cycle of a second-harmonic generation (SHG) cavity for UV light by using an FPGA-based fast control to length-stabilize the cavity for a short period during the experiment only and move it off-resonant otherwise. We demonstrate an application of our technique by carrying out detailed spectroscopy on the isotope shift of the diatomic molecule AlCl, as described in Section 2.2. We note that recently a similar technique that reduces the duty cycle where the circulating power inside a SHG cavity was quickly switched ($\approx\mu$s) with an acousto-optic modulator has been developed [28]. Both of these methods are particularly suitable for experiments geared towards fundamental research since the lack of readily available solid-state UV technology often requires the use of SHG cavities to produce UV light.
@@ 2 Experimental Method
The experiment starts with a cryogenic helium buffer-gas beam source [57] to produce a beam of aluminum monochloride (AlCl) molecules. AlCl molecules are brought into the gas phase via short-pulsed laser ablation of a solid precursor made of a mixture of Al and KCl [58] at 532 nm with a Nd:Yag laser. The molecular beam travels through our vacuum system and is subject to laser-induced fluorescence spectroscopy $\approx 0.5$ m downstream from the source. The time-resolved fluorescence is collected with a photomultiplier tube (Hamamatsu, H10722-04) and analyzed. In the following, we focus the discussion on the spectroscopy laser setup, while further details on the cryogenic buffer-gas beam part of our apparatus can be found in Refs. [23, 24].
@@ 2.1 Resonance-Triggered Stabilization Technique
The spectroscopy light at $261.5$ nm is produced by frequency-doubling CW laser light at 523 nm (Vexlum, VALO SHG SF) in a homebuilt second-harmonic generation (SHG) bow-tie cavity [59] with a non-linear beta-barium-borate (BBO) crystal (Newlight Photonics), as shown in Fig. 1. Frequency drifts of the Vexlum laser are compensated by comparing it to a wavelength meter (High Finesse, WS-7). Specifically, a part of the light from a monitoring output of the laser at 1046 nm is continuously measured by the wavelength meter and compared to a set frequency to generate an error signal. This signal is then fed into a software-based proportional-integral controller (PI) to produce a control signal, which is sent to the internal frequency-tuning piezo of the laser via the analog voltage output of a microcontroller (Arduino Due).
@@
The output of the 523 nm laser is sent through an optical isolator (Thorlabs, IO-5-532-HP) and focused into the SHG cavity using mode-shaping lenses to hit the optimal target beam waist of $\approx 20\,\mu$m inside the crystal, which is determined by the Boyd Kleinman theory [61, 56]. For the SHG process in the BBO at this wavelength, critical type I phase matching at room temperature is achieved by using a crystal that is cut at an angle of $\theta=48.9$ deg. The crystal facets are cut at the Brewster angle for the fundamental wavelength ($\theta\approx 59.2$ deg) to minimize reflection losses at each round trip in the cavity. The cavity is enclosed in a box made of acrylic, as shown in Fig. 5, and the box is continuously flushed with compressed air from the building supply lines. To minimize contamination, an air filter that blocks particulates larger than 0.01 µm (Parker, 9933-11-BQ) is installed inline. After flushing the enclosure for $\approx 10$ min, a humidity level of $\leq 10\%$ and a clean environment is achieved around the cavity. The measured finesse of the cavity is approximately $140$ and the incoupling efficiency is $\approx 27\%$. For the measurements presented in this work, we typically operated with a few milliwatt UV power. We didn’t optimize these cavity parameters further due to the degradation issues of the crystal. Moreover, the measurements presented here only needed low UV power but required us to still use our resonance-triggered stabilization technique. For future experiments that require higher UV power, we will use a CLBO crystal for the SHG cavity, as described in Section 2.3.
@@
The complete experimental sequence, shown in Fig. 2, is as follows: First, the main laser wavelength is tuned and stabilized to a given set point by the computer. Then, our data acquisition system Artiq [68] sends a TTL trigger (T) to the FPGA board to initiate the voltage ramp (Ramp). The FPGA switches on the cavity stabilization (PID on) as soon as it reaches the resonance condition. After a short delay, which is chosen such that the cavity resonance condition is always met by ramping over at least one free spectral range, Artiq triggers an ablation laser shot (Yag) to produce a beam of molecules. The SHG cavity is then held on resonance for a time interval that is sufficient to let the molecules traverse the experimental apparatus ($\approx 50$ ms) and be detected on a photomultiplier (PMT detection). After this time, the PI control module is switched off and the piezo voltage ramp output is set to zero to intentionally tune the cavity out of resonance and eliminate the UV output. This sequence is repeated for a preset number of averages after which the laser wavelength is tuned to a new set point. An example of the voltage ramp (solid blue curve) and the UV output intensity (solid orange curve) of a single experimental cycle is shown in Fig. 2.
@@ 2.2 Measurement and Ab Initio Calculations of the Isotope Shift in AlCl
Using the method described in Section 2.1, we carried out detailed spectroscopy on the isotope shift of the $X^{1}\Sigma(v=0)\leftarrow A^{1}\Pi(v^{\prime}=0)$ $Q$-transitions at 261.5 nm in AlCl. Fig. 3 shows the laser-induced fluorescence spectrum of both isotopologues, Al35Cl and Al37Cl, as a function of the laser detuning. The blue dots are the measured data and each point corresponds to 30 averages, and the red solid curve corresponds to a Gaussian multi-peak fit to identify the center of each line. This data was taken using the calibrated wavelength meter, as described above. The difference in the amplitudes of two isotope manifolds comes from the natural abundances of 75.8% for 35Cl and 24.2% for 37Cl. The individual peaks of each manifold are a combination of the $Q(1),Q(2),\ldots$-transitions and the hyperfine structure splittings of the involved states that originates from the nuclear spins of Al ($I_{1}=5/2$) and 35/37Cl ($I_{2}=3/2$) [23, 24]. The short lifetime of $6$ ns of the $A^{1}\Pi$ electronic state results in a broad natural linewidth ($\approx 27$ MHz) that does not resolve the chlorine hyperfine splitting which spreads over $\approx 11.5$ MHz [69]. The similarity of the rotational constants of the $X^{1}\Sigma$- and $A^{1}\Pi$-states [23] results in an overlap of the first few $Q$ transitions. Therefore, we chose to compare the frequency of the fitted combined peaks, labeled $a,b,\dots,g$ and $a^{\prime},b^{\prime},\dots,g^{\prime}$, to extract an average isotope shift. The result of this comparison is shown in the inset of Fig. 3 for all seven peaks with an average of $6124\pm 8$ MHz. This measurement is consistent with the isotope shift extracted from the lower resolution absorption spectroscopy in our previous work, which yields $\approx 6128$ MHz [23] and is the first report of the electronic isotope shift in AlCl to the best of our knowledge.
@@
The average theoretical isotope shift is $5881$ MHz which is within $4\%$ of the experimental value. This level of agreement is considered excellent. The small differences are attributed to the residual numerical errors in the ab initio calculations, potential energy curve fitting errors, and numerical errors in the rovibrational calculations. We note that the $c,c^{\prime}$ and $e,e^{\prime}$ features observed in the experimental spectra are most likely associated with hyperfine structure within the overlapping $Q$ transitions and is not included in the current level of theory. The theoretical frequencies listed in Table 1 are used to simulate the experimental spectra by adding up the appropriately weighted contributions from each $Q$ transition with a Lorentzian line-width function [71] centered at each transition frequency. The line widths were chosen to increase linearly with the rotational quantum number $J$ via $15\,(J+1)$ MHz to qualitatively reproduce the experimentally observed widths of the $Q$ peaks in Fig. 3. The weights of each $Q$ transition were chosen to match the relative intensities of the experimental spectra for Al35Cl in Fig. 3. The weights for the Al37Cl $Q$ transitions were set by scaling the Al35Cl weights by the relative isotope abundance ratio: $0.24/0.76=0.32$.
Overall, the theoretically simulated spectra presented in Fig. 3 reproduces the experimental one quite well confirming the assignments of the primary overlapping $Q$ branch features. The theoretical analysis also confirms that the experimentally measured isotope shift is due to an increase in the relative difference between the vibrational zero point energies (ZPE) for the $X$ and $A$-states upon isotopic substitution. That is, the vibrational ZPEs for the $X$ and $A$ states of Al35Cl are $240.047\,{\rm cm}^{-1}$ and $224.751\,{\rm cm}^{-1}$, respectively. Whereas for Al37Cl the ZPEs for the $X$ and $A$ states are $237.213\,{\rm cm}^{-1}$ and $222.108\,{\rm cm}^{-1}$. The corresponding isotope shifts in the vibrational ZPEs for the $X$ and $A$ states are therefore $-2.834,{\rm cm}^{-1}$ and $-2.643,{\rm cm}^{-1}$. Thus, we see that the ZPE for the $X$ state shifts lower in energy than that for the $A$ state and the relative difference increases by $0.191\,{\rm cm}^{-1}$ (or $5726$ MHz). The increased relative difference in vibrational ZPE accounts for $97\%$ of the observed isotope shift ($5726/5881=0.97$). The remaining $3\%$ is due to shifts in the rotational energies due to changes in the rotational constants upon isotopic substitution. We note that the difference in the vibrational ZPE shifts between the $A$ and $X$ states is due to the difference in the potential energy curves (e.g., force constants). Also, the electronic transition energy $T_{e}$ is not affected by isotopic substitution and is the same for both isotopologues.

### v2

# Low Duty Cycle Pulsed UV Technique for Spectroscopy of Aluminum Monochloride

@@
Low Duty Cycle Pulsed UV Technique for Spectroscopy of Aluminum Monochloride
@@ 1 Introduction
The invention of the laser led to a myriad of applications in science and industry [1]. In fundamental research, the laser enabled applications ranging from precision spectroscopy and tests of fundamental theories [2], cooling and trapping of atoms and molecules [3, 4], to cold controlled chemistry [5, 6, 7]. This type of research often uses narrow-band continuous-wave (CW) lasers to provide a high frequency resolution. Many of these applications also require laser light in the deep ultraviolet (UV) to cover optical transitions in atoms and molecules for carrying out precision spectroscopy and laser cooling [3]. Specific examples include atomic cadmium with a laser cooling transition at 229 nm [8, 9, 10] and atomic mercury at 254 nm [11, 12].
@@
In addition to a specific wavelength in the deep-UV, many of the interesting applications require the use of high UV laser power. For instance, driving dipole-forbidden transitions, as is done for spectroscopy of hydrogen [15, 16, 17, 18], muonium [19] and xenon [20], requires deep-UV laser light at the level of Watts. Another example involves the laser cooling of molecules [21, 4, 22, 23]. Two of the promising candidates, AlCl [24, 25] and AlF [26], have laser cooling transitions at 261 nm and 228 nm, respectively. These molecules potentially provide high capture velocities for magneto-optical traps of up to 30–40 m/s [25]. To achieve this level, though, high intensities are required to saturate the cooling transitions. At the same time, this requirement is in conflict with the desire to use laser beams with large beam cross sections. Here, the aim is to provide a large overlap with a molecular beam to render the capture process as efficient as possible. To address such experimental needs, tremendous progress has been made over the years and various laser systems at wavelengths in the deep-UV below 300 nm with output levels ranging from 50 mW to more than 1 Watt, have been developed [27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 10].
@@
A major challenge when using deep-UV lasers is the laser-induced damage that can occur in any of the involved optical components. The intensity threshold, typically designated as laser-induced damage threshold (LIDT), at which significant damage occurs depends on the optical material itself and varies with the wavelength. The mechanisms of these effects have been studied for typical used optical materials, e.g. CaF2 [39, 40] and fused silica [41, 42, 43]. Among the detrimental effects causing damage is the absorption of moisture by materials that are hygroscopic, the breaking of chemical bonds, the depletion of oxygen and the contamination with hydrocarbons upon exposure [44, 45, 46]. Various methods have been explored to revert these effects that focus on keeping oxygen near the surface of the optics through oxide coatings [46] or the submersion of the optics in an oxygen environment [33]. Moreover, fluoride-coated optics that avoid the need for the presence of oxygen have been shown to be beneficial [47]. Non-linear crystals, which are used for higher-harmonic generation to produce deep-UV light, are often hygroscopic and suffer from UV-induced damage. Extensive studies have been carried out to characterize these damages and to find mitigation techniques for non-linear crystals, such as BaB2O4 (BBO) [48, 49, 50] and CsLiB6O10 (CLBO) [51, 52, 53, 54, 50], which are relevant for this work. Optical fibers, another essential component in today’s experiments involving lasers, have been shown to be able to withstand larger amounts of UV intensities if they are pretreated with hydrogen and irradiate with UV light [55].
@@
While the choice of materials and the developed specific treatments have shown promising results to prepare optical components for operating in the deep-UV regime, a material-independent method to reduce UV-degradation in the first place is to minimize high intensity UV-exposure of an optical material. In some optical setups this is possible by avoiding laser beam foci near or inside optical materials, which consequently reduces the local UV intensity [56]. This technique, however, cannot be used in setups where optical foci are a requirement. Here, an alternative approach is to reduce the duty cycle of the UV light to a minimum instead. This technique is most naturally applied in pulsed experiments that often exhibit times between experimental repeats where the laser is only idling. However, it can also be used to bridge dead times in continuous experiments.
@@
In this work, we present a method to reduce the duty cycle of a second-harmonic generation (SHG) cavity for UV light by using an FPGA-based fast control to length-stabilize the cavity for a short period during the experiment only and move it off-resonant otherwise. We demonstrate an application of our technique by carrying out detailed spectroscopy on the isotope shift of the diatomic molecule AlCl, as described in Section 2.2. We note that recently a similar technique that reduces the duty cycle where the circulating power inside a SHG cavity was quickly switched ($\approx\mu$s) with an acousto-optic modulator has been developed [29]. Both of these methods are particularly suitable for experiments geared towards fundamental research since the lack of readily available solid-state UV technology often requires the use of SHG cavities to produce UV light. As a consequence of the reduced duty cycle, any optics that is used to steer or manipulate the UV laser light for the experiment experiences significantly less UV radiation exposure. This, in turn, prolongs the lifetime of the optical setups and renders experiments that use deep-UV light more cost-effective.
In the future, we plan to use our pulsed approach to implement a robust, high-power laser system to cool and load AlCl in a magneto-optical trap. Laser cooling of AlCl is expected to produce a large number of trapped molecules due to the high capture velocity, but requires high laser powers. The large trapped numbers are beneficial to many proposed experiments with cold molecules, including controlled chemistry [5, 57, 58, 59], quantum simulation [60, 61, 62, 6, 63, 64, 65] or precision measurements [66, 67, 68, 69, 70, 71, 72, 73, 67, 74, 75, 76, 77, 78, 79, 80, 81, 82, 83, 84, 85, 86].
@@ 2 Experimental Method
The experiment starts with a cryogenic helium buffer-gas beam source [87] to produce a beam of aluminum monochloride (AlCl) molecules. AlCl molecules are brought into the gas phase via short-pulsed laser ablation of a solid precursor made of a mixture of Al and KCl [88] at 532 nm with a Nd:Yag laser. The molecular beam travels through our vacuum system and is subject to laser-induced time-resolved fluorescence spectroscopy $\approx 0.5$ m downstream from the source. To provide a good signal-to-noise, the fluorescence light is imaged onto a relay image using a plano-convex lens (focal length of 75 mm) and spatially filtered with an iris. The relay image is then focused with two plano-convex lenses (focal length of 100 mm) onto a photo multiplier tube (Hamamatsu, H10722-04) and analyzed. A UV bandpass filter (Thorlabs, FGUV5M) with a transmission of 70% at 260 nm is used to remove background scatter onto the PMT. In the following, we focus the discussion on the spectroscopy laser setup, while further details on the cryogenic buffer-gas beam part of our apparatus can be found in Refs. [24, 25].
@@ 2.1 Resonance-Triggered Stabilization Technique
The spectroscopy light at $261.5$ nm is produced by frequency-doubling CW laser light at 523 nm (Vexlum, VALO SHG SF) in a homebuilt second-harmonic generation (SHG) bow-tie cavity with a non-linear beta-barium-borate (BBO) crystal (Newlight Photonics), as shown in Fig. 1. The cavity consists of two curved mirrors with a radius of curvature of $50$ mm, an input coupler with 97% transmission at 523 nm and a piezo-actuated mirror for length stabilization. Frequency drifts of the Vexlum laser are compensated by comparing it to a wavelength meter (High Finesse, WS-7). Specifically, a part of the light from a monitoring output of the laser at 1046 nm is continuously measured by the wavelength meter and compared to a set frequency to generate an error signal. This signal is then fed into a software-based proportional-integral controller (PI) to produce a control signal, which is sent to the internal frequency-tuning piezo of the laser via the analog voltage output of a microcontroller (Arduino Due).
@@
The output of the 523 nm laser is sent through an optical isolator (Thorlabs, IO-5-532-HP) and focused into the SHG cavity using mode-shaping lenses to hit the optimal target beam waist of $\approx 20\,\mu$m inside the crystal, which is determined by the Boyd Kleinman theory [90, 91]. For the SHG process in the BBO at this wavelength, critical type I phase matching at room temperature is achieved by using a crystal that is cut at an angle of $\theta=48.9$ deg. The crystal facets are cut at the Brewster angle for the fundamental wavelength ($\theta\approx 59.2$ deg) to minimize reflection losses at each round trip in the cavity. The cavity is enclosed in a box made of acrylic with dimensions 62 x 33 x 19 cm , as shown in Fig. 2, and the box is continuously flushed with compressed air from the building supply lines with an airflow of $\approx 1$ liter/min. To minimize contamination, an air filter that blocks particulates larger than 0.01 µm (Parker, 9933-11-BQ) is installed inline. After flushing the enclosure for $\approx 10$ min, a humidity level of $\leq 10\%$ and a clean environment is achieved around the cavity. The measured finesse of the cavity is approximately $140$ and the incoupling efficiency is $\approx 27\%$. For the measurements presented in this work, we typically operated with a few milliwatt UV power. We didn’t optimize these cavity parameters further due to degradation issues of the crystal. Moreover, the measurements presented here only needed low UV power but required us to still use our resonance-triggered stabilization technique. For future experiments that require higher UV power, we will use a CLBO crystal for the SHG cavity. More details on these points are discussed in Section 2.3.
@@
The complete experimental sequence, shown in Fig. 3, is as follows: First, the main laser wavelength is tuned and stabilized to a given set point by the computer. Then, our data acquisition system Artiq [99] sends a TTL trigger (T) to the FPGA board to initiate the voltage ramp (Ramp). The FPGA switches on the cavity stabilization (PID on) as soon as it reaches the resonance condition. After a short delay, which is chosen such that the cavity resonance condition is always met by ramping over at least one free spectral range, Artiq triggers an ablation laser shot (Yag) to produce a beam of molecules. The SHG cavity is then held on resonance for a time interval that is sufficient to let the molecules traverse the experimental apparatus ($\approx 50$ ms) and be detected on the photomultiplier (PMT detection). After this time, the PI control module is switched off and the piezo voltage ramp output is set to zero to intentionally tune the cavity out of resonance and eliminate the UV output. This sequence is repeated for a preset number of averages after which the laser wavelength is tuned to a new set point. An example of the voltage ramp (solid blue curve) and the UV output intensity (solid orange curve) of a single experimental cycle is shown in Fig. 3.
@@ 2.2 Measurement and Ab Initio Calculations of the Isotope Shift in AlCl
Using the method described in Section 2.1, we carried out detailed spectroscopy on the isotope shift of the $X^{1}\Sigma(v=0)\leftarrow A^{1}\Pi(v^{\prime}=0)$ $Q$-transitions at 261.5 nm in AlCl. Fig. 4 shows the laser-induced fluorescence spectrum of both isotopologues, Al35Cl and Al37Cl, as a function of the laser detuning. The blue dots are the measured data and each point corresponds to 30 averages, and the red solid curve corresponds to a Gaussian multi-peak fit to identify the center of each line. This data was taken using the calibrated wavelength meter, as described above. The difference in the amplitudes of two isotope manifolds comes from the natural abundances of 75.8% for 35Cl and 24.2% for 37Cl. The individual peaks of each manifold are a combination of the $Q(1),Q(2),\ldots$-transitions and the hyperfine structure splittings of the involved states that originates from the nuclear spins of Al ($I_{1}=5/2$) and 35/37Cl ($I_{2}=3/2$) [24, 25]. The short lifetime of $6$ ns of the $A^{1}\Pi$ electronic state results in a broad natural linewidth ($\approx 27$ MHz) that does not resolve the chlorine hyperfine splitting which spreads over $\approx 11.5$ MHz [100]. The similarity of the rotational constants of the $X^{1}\Sigma$- and $A^{1}\Pi$-states [24] results in an overlap of the first few $Q$ transitions. Therefore, we chose to compare the frequency of the fitted combined peaks, labeled $a,b,\dots,g$ and $a^{\prime},b^{\prime},\dots,g^{\prime}$, to extract an average isotope shift. The result of this comparison is shown in the inset of Fig. 4 for all seven peaks with an average of $6124\pm 8$ MHz. This measurement is consistent with the isotope shift extracted from the lower resolution absorption spectroscopy in our previous work, which yielded $\approx 6128$ MHz [24] and is the first report of the electronic isotope shift in AlCl of the $X(v=0)\leftarrow A(v^{\prime}=0)$ $Q$-transition. To the best of our knowledge, the only available other detailed work on isotope shifts in AlCl is Nakadi et al. [101]. Due to their instrument’s resolution of $\approx 1.5$ pm, they were not able to observe this transition and reported measurements only for higher vibrational states ($v\geq 1$). The characterization of the $X(v=0)\leftarrow A(v^{\prime}=0)$ transition, however, is important for future atomic physics experiments with AlCl since it is part of the laser cooling scheme of AlCl.
@@
The differences between the experimental and theoretical Q line centers are quite reasonable considering that no adjustable parameters were optimized to fit the theoretically computed rotational spectra to the experimentally measured one [25]. The average theoretical isotope shift is $5881$ MHz which is within $4$% of the experimental value. This level of agreement is considered excellent for ab initio based theory with no adjustable parameters [25]. We note that the $c,c^{\prime}$ and $e,e^{\prime}$ features observed in the experimental spectra are most likely associated with hyperfine structure within the overlapping $Q$ transitions and is not included in the current level of theory. The theoretical frequencies listed in Table 1 are used to simulate the experimental spectra by adding up the appropriately weighted contributions from each $Q$ transition with a Lorentzian line-width function [103] centered at each transition frequency. The line widths were chosen to increase linearly with the rotational quantum number $J$ via $15\,(J+1)$ MHz to qualitatively reproduce the experimentally observed widths of the $Q$ peaks in Fig. 4. The weights of each $Q$ transition were chosen to match the relative intensities of the experimental spectra for Al35Cl in Fig. 4. The weights for the Al37Cl $Q$ transitions were set by scaling the Al35Cl weights by the relative isotope abundance ratio: $0.24/0.76=0.32$.
Overall, the theoretically simulated spectra presented in Fig. 4 reproduces the experimental one quite well confirming the assignments of the primary overlapping $Q$ branch features. The theoretical analysis also confirms that the experimentally measured isotope shift is due to an increase in the relative difference between the vibrational zero point energies (ZPE) for the $X$ and $A$-states upon isotopic substitution. That is, the vibrational ZPEs for the $X$ and $A$ states of Al35Cl are $240.047\,{\rm cm}^{-1}$ and $224.751\,{\rm cm}^{-1}$, respectively. Whereas for Al37Cl the ZPEs for the $X$ and $A$ states are $237.213\,{\rm cm}^{-1}$ and $222.108\,{\rm cm}^{-1}$. The corresponding isotope shifts in the vibrational ZPEs for the $X$ and $A$ states are therefore $-2.834\,{\rm cm}^{-1}$ and $-2.643\,{\rm cm}^{-1}$. Thus, we see that the ZPE for the $X$ state shifts lower in energy than that for the $A$ state and the relative difference increases by $0.191\,{\rm cm}^{-1}$ (or $5726$ MHz). The increased relative difference in vibrational ZPE accounts for $97\%$ of the observed isotope shift ($5726/5881=0.97$). The remaining $3\%$ is due to shifts in the rotational energies due to changes in the rotational constants upon isotopic substitution. We note that the difference in the vibrational ZPE shifts between the $A$ and $X$ states is due to the difference in the potential energy curves (e.g., force constants). Also, the electronic transition energy $T_{e}$ is not affected by isotopic substitution and is the same for both isotopologues.